40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 4 02:59:05 2014, ericq, Update, LSC, Bump in Darm Plant Dec4_darmbump.pdf
    Reply  Thu Dec 4 03:09:06 2014, Jenne, Update, LSC, Bump in Darm Plant: Lockloss after measurement 8x
Message ID: 10754     Entry time: Thu Dec 4 02:59:05 2014     Reply to this: 10755
Author: ericq 
Type: Update 
Category: LSC 
Subject: Bump in Darm Plant 

[J,Q]

After some housekeeping (ASS is wonky, alignment of X green beat was bad, tuning of demod angles, fm gains for REFL165), we were able to bring the PRFPMI up to arm powers of 8 very stably. 

We were keeping an eye on the DARM OLG, to make sure the gain was correct. We then saw a bump around 120Hz. Here is the bump. 

Dec4_darmbump.pdf

Changing CARM offset changes its amplitude. Maybe it's a DARM optical spring. It didn't occur to me until after we lost lock that we could have tweaked the DARM offset to move it around if this was the case. 

Unfortunately, due to some unexplained locklosses, we weren't able to get back into a state to measure this more... which is annoying. During that stable lock, Jenne stated that PRCL and DARM noises were looking particularly good. 

We may want to tweak the way we handle the transmission PD handoff; maybe we want to force the switch at a certain place in the carm_up script, so that we're not flipping back and forth during an IR handoff; I think this may have been responsible for a lock loss or two. 

ELOG V3.1.3-